# Resilient Topology-Aware Coordination for Dynamic 3D UAV Networks under Node Failure

### v1

@@ 5.1. Overview of the TAG-MAPPO Architecture
5.2. Dynamic Ego-Centric Graph Representation
@@
To formalize the time-varying interaction structure, the observation of the $k$-th UAV is modeled as a dynamic ego-centric graph $\mathcal{G}_{k}(t)=(\mathcal{V}_{k}(t),\mathcal{E}_{k}(t))$. The node set $\mathcal{V}_{k}(t)$ incorporates the ego-UAV itself, the fixed Ground Base Station (GBS), and a dynamic set of neighboring UAV agents $\mathcal{N}_{k}(t)$ located within the communication range $R_{\mathrm{comm}}$. Crucially, in high-mobility scenarios, the cardinality of the neighbor set $|\mathcal{N}_{k}(t)|$ is highly volatile. As UAVs and users adjust their positions or when a node failure occurs, topological links are frequently established or broken. Standard MLPs struggle to process such variable-length inputs efficiently because they often require zero-padding that introduces computational waste or numerical noise. In contrast, our graph-based formulation allows the critic to naturally handle these topological fluctuations without structural modification. The attention mechanism can autonomously redistribute priority weights or ignore the absence of failed nodes, ensuring stable value estimation across fragmented network states.
5.3. Feature Extraction and Embedding with Shuffling
@@
Before applying the attention mechanism, raw sensory data must be transformed into high-level latent representations. Reflecting the CTDE architecture, distinct encoders are employed for the decentralized actors and the centralized critic to cater to their specific functional roles.
@@
For each decentralized actor, the local raw observation is mapped into a compact local embedding vector via a lightweight MLP encoder. This encoder is shared across all UAV agents to facilitate parameter efficiency and cooperative behavior. The resulting local embedding encapsulates the immediate kinematic state required for reactive control and serves as the primary input for the policy head.
@@
The centralized critic employs a more sophisticated encoding scheme to handle the global state information. To ensure that the learned policy remains robust against the indexing order of agents, we introduce a Random Observation Shuffling (ROS) mechanism during the training phase. Before entering the TA-GAT layer, the neighbor feature set is processed by a stochastic permutation operator to achieve Permutation Invariance:
@@
(20) $\tilde{\mathbf{X}}_{k}^{\text{neigh}}(t)=\text{StochasticPermutation}\left(\{\mathbf{h}_{j}(t)\}_{j\in\mathcal{N}_{k}(t)}\right).$
This ensures that the critic learns to prioritize neighbors based on their physical attributes rather than their position in the input tensor. By projecting these diverse inputs into a unified feature space, the encoders enable the TA-GAT layer to perform effective relational reasoning under varying topological densities.
@@
Crucially, the ROS mechanism is applied only to the neighboring agent set $\{\mathbf{h}_{j}(t)\}_{j\in\mathcal{N}_{k}(t)}$, while the ego-state feature $\mathbf{h}_{k}(t)$ remains fixed as an anchor for the attention query. This decoupled shuffling strategy forces the attention mechanism to learn the relative topological relationships between the target UAV and its surrounding swarm. By treating the neighborhood as an unordered collection of relational entities, the policy becomes resilient to dynamic changes in the agent population.
@@
This design directly contributes to the system’s self-healing capabilities. When a neighboring node disappears due to hardware failure, the remaining agents can maintain stable coordination because the learned feature aggregation is independent of specific neighbor indices. By projecting these diverse and permuted inputs into a unified feature space, the encoders enable the TA-GAT layer to perform effective relational reasoning under varying topological densities and volatile operational conditions.
5.4. Topology-Aware Graph Attention Mechanism (TA-GAT)
@@
A fundamental challenge in applying Graph Neural Networks to UAV control is the potential dilution of the agent’s own state information when aggregating neighbor features. To address the challenge of topology deformation, we introduce a Dual-Path Aggregation strategy within the TA-GAT layer.
5.4.1. Attention-based Neighbor Aggregation
@@
(21) $e_{kj}(t)=\frac{(\mathbf{W}_{Q}\mathbf{h}_{k}(t))^{T}(\mathbf{W}_{K}\mathbf{h}_{j}(t))}{\sqrt{d_{\text{attn}}}}.$
@@
(22) $\mathbf{c}_{\text{neigh}}(t)=\sum_{j\in\mathcal{V}(t)\setminus\{k\}}\alpha_{kj}(t)(\mathbf{W}_{V}\mathbf{h}_{j}(t)).$
@@
5.4.2. Dual-Path Fusion and Skip Connection
@@
(23) $\mathbf{g}_{k}(t)=\text{MLP}_{\text{head}}\left([\mathbf{W}_{\text{self}}\mathbf{h}_{k}(t)\mathbin{\|}\mathbf{c}_{\text{neigh}}(t)]\right),$
@@
5.5. Multi-Agent Actor-Critic Learning
@@
(24) $L(\theta)=\mathbb{E}_{k,t}\left[\min(\rho_{k}(t)A_{k}(t),\text{clip}(\rho_{k}(t),1-\epsilon,1+\epsilon)A_{k}(t))\right],$
@@
(25) $\mathcal{L}(\phi)=\begin{cases}\frac{1}{2}(y_{t}-V_{\phi}(s_{t}))^{2}&\text{for }|y_{t}-V_{\phi}(s_{t})|\leq\delta,\\ \delta(|y_{t}-V_{\phi}(s_{t})|-\frac{1}{2}\delta)&\text{otherwise},\end{cases}$
@@
5.6. TAG-MAPPO Training Algorithm
The complete training procedure for the TAG-MAPPO framework is formalized in Algorithm 1. The optimization pipeline is structured into two distinct phases, namely decentralized data collection and centralized parameter updates.
@@ 5.6. TAG-MAPPO Training Algorithm
In the first phase, distributed UAV agents interact with the environment using their lightweight actor networks. At each time step, agents rely solely on local observations to sample actions, which ensures that the inference latency remains constant regardless of the network size. If a node failure occurs during this phase, the surviving swarm $\mathcal{V}(t)$ must immediately adapt its coordination based on the remaining local observations. The transition data, including the instantaneous topology state $\mathcal{V}(t)$, is stored in the replay buffer to provide the critic with the necessary context for post-failure value estimation.
@@
In the second phase, the system executes centralized updates. Before optimization, the Generalized Advantage Estimation (GAE) is computed to balance the bias-variance trade-off. During this stage, the TA-GAT critic plays a pivotal role by reconstructing the dynamic ego-centric graph for each sample in the minibatch. Specifically, the ROS mechanism is applied to the neighbor feature set before graph aggregation. By stochastically reordering the input neighbors, the critic is forced to learn relational features that are invariant to agent indexing. By executing the dual-path aggregation, the critic predicts state values to compute the Huber-based value loss. This mechanism ensures that the lightweight actors are trained to be topology-aware by distilling the complex relational knowledge of the critic into the simplistic policy networks.
5.7. Computational Complexity Analysis
@@
During the online execution phase, the computational complexity is dominated by matrix-vector multiplications within the MLP-based actors, which scales as $\mathcal{O}(L\cdot D^{2})$ where $L$ is the number of layers and $D$ is the hidden dimension. This complexity is strictly independent of the number of dynamic neighbors, which satisfies the real-time constraints required for 6G mission-critical tasks.
@@
During the offline training phase, the complexity is primarily driven by the attention mechanism within the centralized critic. For each agent, the TA-GAT module involves linear projections and scaled dot-product computations. For an ego-centric graph, this scales as $\mathcal{O}((|\mathcal{V}(t)|+|\mathcal{N}_{k}(t)|)\cdot D)$, where $|\mathcal{N}_{k}(t)|$ is the number of active neighbors. While the attention operation involves more intensive processing than simple MLPs, its execution is restricted to the centralized training server with substantial computational resources. Consequently, TAG-MAPPO achieves an optimal trade-off by utilizing topology-aware reasoning offline to distill robust policies into efficient, lightweight actors for deployment.
6. Simulation Results and Analysis
@@
In this section, we conduct extensive simulations to evaluate the performance and resilience of the proposed TAG-MAPPO framework. We first outline the simulation setup and then analyze the experimental results across various environmental scenarios, with a particular focus on the system recovery capability following node failures.
6.1. Simulation Setup and Parameters
@@
The simulation environment is developed using a customized Python-based multi-UAV simulator. The target area is defined as a 1000×1000 $\times 1000\text{ m}^{2}$ square region served by a heterogeneous network consisting of $K_{U}=4$ mobile UAVs and one stationary ground base station (GBS), totaling $K=5$ serving nodes for $M=140$ ground users. The flight altitude corridor for UAVs is set between $H_{\min}=80\text{ m}$ and $H_{\max}=120\text{ m}$. To accurately reflect heterogeneous mobility, the maximum UAV flight speed is capped at $V_{\max}^{\text{UAV}}=30\text{ m/s}$, while the maximum speed for independent ground users is $V_{\max}^{\text{user}}=15\text{ m/s}$. In clustered scenarios, the group velocity for vehicular platoons is maintained at $10\text{ m/s}$.
@@
To ensure balanced learning across multiple objectives, all individual reward components in (14) are normalized to the range of $[0,1]$. The weighting coefficients are empirically set as $(\lambda_{\text{Cov}},\lambda_{\text{EE}},\lambda_{\text{JR}},\lambda_{\text{JL}},\lambda_{\text{Min}})=(2.0,0.1,0.5,0.8,0.5)$, which prioritizes global coverage while maintaining service quality for the worst-case user. Additionally, the handoff penalty coefficient $\lambda_{\text{HO}}$ is specifically tuned to $0.1$ to suppress redundant signaling and maintain topological stability during reconfiguration phases.
@@
To evaluate system resilience, a catastrophic node failure is programmed at $t=100$ by setting the survival indicator $\xi_{k}(t)=0$ for a randomly selected UAV node (excluding the GBS), forcing the remaining swarm to reconfigure the network topology autonomously. The consolidated parameters are detailed in Table 1.
Table 1. Consolidated Simulation Parameters and Hyperparameter Settings
Parameter (Physical) Value Hyperparameter (DRL) Value
Number of UAVs ($K_{U}$) 4 Actor Learning Rate $1\times 10^{-4}$
Ground Base Station (GBS) 1 Critic Learning Rate $5\times 10^{-4}$
UAV Max Speed ($V_{\max}^{\text{UAV}}$) 30 m/s Discount Factor ($\gamma$) 0.99
User Max Speed ($V_{\max}^{\text{user}}$) 15 m/s GAE Parameter ($\tau$) 0.95
Carrier Frequency ($f_{c}$) 2.0 GHz PPO Epochs 10
System Bandwidth ($B$) 20 MHz Batch Size 256
Transmission Power ($P_{\text{tx}}$) 23 dBm Huber Loss Delta ($\delta$) 2.0
Noise PSD ($N_{0}$) 174-174 dBm/Hz Entropy Coeff. ($\beta_{\text{ent}}$) $0.10\to 0.01$
Power & Geometry Value Reward Weights Value
Altitude Range $H_{\min},H_{\max}] [$80,120][80,120] m Coverage ($\lambda_{\text{Cov}}$) 2.0
GBS Fixed Power ($P_{\text{GBS}}$) 20 W Energy Efficiency ($\lambda_{\text{EE}}$) 0.1
UAV Comm. Power ($P_{\text{comm}}$) 5 W Rate Fairness ($\lambda_{\text{JR}}$) 0.5
Target Area $\mathcal{D}$ 1×1 $\times 1\text{ km}^{2}$ Load Fairness ($\lambda_{\text{JL}}$) 0.8
Safety Distance ($d_{\text{safe}}$) 10 m Minimum Rate ($\lambda_{\text{Min}}$) 0.5
Number of Users ($M$) 140 Handoff Penalty ($\lambda_{\text{HO}}$) 0.1
6.2. Benchmark Scenarios and Channel Characterization
@@
To evaluate the generalization and robustness of the TAG-MAPPO framework, we define three distinct simulation scenarios based on the physical parameters in the 3GPP propagation models. The specific LoS/NLoS characteristics for each environment are detailed as follows:
@@
  • Crowded Urban: This scenario represents an interference-limited regime with high user density. We set the channel parameters as $(a=12.08,b=0.11)$ with a significant NLoS path loss exponent ($\eta_{\text{nlos}}=23.0$). The primary challenge is managing complex co-channel interference within user clusters modeled by the RPGM mobility.
@@
  • Suburban: This scenario features moderate user dispersion and a more balanced LoS probability ($a=9.61,b=0.16$). It serves as the baseline for evaluating the ”optimal resilience” of the UAV swarm, as the topology allows for effective spatial reuse.
@@
  • Rural Area: Characterized by sparse, Gauss-Markov independent mobility, this scenario is a noise-limited environment with a high LoS probability ($a=4.88,b=0.43$). The lower blockage frequency simplifies individual link optimization but tests the ”over-smoothing” limits of the graph coordination mechanism.
6.3. Performance Metrics

### v2

@@ V-A Overview of the TAG-MAPPO Architecture
V-B Dynamic Ego-Centric Graph Representation
@@
To formalize the time-varying interaction structure, the observation of the $k$-th UAV is modeled as a dynamic ego-centric graph $\mathcal{G}_{k}(t)=(\mathcal{V}_{k}(t),\mathcal{E}_{k}(t))$. The node set $\mathcal{V}_{k}(t)$ incorporates the ego-UAV, the fixed GBS, and a dynamic set of neighboring UAV agents $\mathcal{N}_{k}(t)$ within communication range $R_{\mathrm{comm}}$. In high-mobility scenarios, the cardinality of the neighbor set $|\mathcal{N}_{k}(t)|$ is volatile. As UAVs and users adjust positions or when node failures occur, topological links are frequently established or broken. Standard MLPs struggle to process such variable-length inputs because they often require zero-padding, which introduces computational waste or numerical noise. In contrast, our graph-based formulation allows the critic to handle these topological fluctuations naturally. The attention mechanism can autonomously redistribute priority weights or ignore the absence of failed nodes, ensuring stable value estimation across fragmented network states.
V-C Feature Extraction and Embedding with Shuffling
@@
Before applying the attention mechanism, raw sensory data are transformed into high-level latent representations. Reflecting the CTDE architecture, distinct encoders are employed for decentralized actors and the centralized critic to cater to their specific functional roles.
@@
For each decentralized actor, the local raw observation is mapped into a compact embedding vector via a lightweight MLP encoder. This encoder is shared across all UAV agents to facilitate parameter efficiency and cooperative behavior. The resulting local embedding encapsulates the immediate kinematic state required for reactive control and serves as the primary input for the policy head.
@@
The centralized critic employs a sophisticated encoding scheme to handle global state information. To ensure the learned policy remains robust against agent indexing, we introduce a Random Observation Shuffling (ROS) mechanism during training. Before entering the TA-GAT layer, the neighbor feature set is processed by a stochastic permutation operator to achieve Permutation Invariance:
@@
$\tilde{\mathbf{X}}_{k}^{\text{neigh}}(t)=\text{StochasticPermutation}\left(\{\mathbf{h}_{j}(t)\}_{j\in\mathcal{N}_{k}(t)}\right).$ (20)
@@
This ensures the critic learns to prioritize neighbors based on physical attributes rather than their position in the input tensor. Crucially, ROS is applied only to the neighboring agent set $\{\mathbf{h}_{j}(t)\}_{j\in\mathcal{N}_{k}(t)}$, while the ego-state feature $\mathbf{h}_{k}(t)$ remains fixed as an anchor for the attention query. This decoupled shuffling strategy forces the attention mechanism to learn relative topological relationships between the target UAV and its surrounding swarm. By treating the neighborhood as an unordered collection of relational entities, the policy becomes resilient to dynamic changes in the agent population.
@@
This design directly contributes to the system’s self-healing capabilities. When a neighboring node fails, remaining agents maintain stable coordination because the learned feature aggregation is independent of specific neighbor indices. By projecting these diverse and permuted inputs into a unified feature space, the encoders enable the TA-GAT layer to perform effective relational reasoning under varying topological densities and volatile operational conditions.
V-D Topology-Aware Graph Attention Mechanism (TA-GAT)
@@
A fundamental challenge in applying Graph Neural Networks to UAV control is the potential dilution of the agent’s own state information during neighbor feature aggregation. To address topology deformation, we introduce a Dual-Path Aggregation strategy within the TA-GAT layer.
V-D1 Attention-based Neighbor Aggregation
@@
$e_{kj}(t)=\frac{(\mathbf{W}_{Q}\mathbf{h}_{k}(t))^{T}(\mathbf{W}_{K}\mathbf{h}_{j}(t))}{\sqrt{d_{\text{attn}}}}.$ (21)
@@
$\mathbf{c}_{\text{neigh}}(t)=\sum_{j\in\mathcal{V}(t)\setminus\{k\}}\alpha_{kj}(t)(\mathbf{W}_{V}\mathbf{h}_{j}(t)).$ (22)
@@
V-D2 Dual-Path Fusion and Skip Connection
@@
$\mathbf{g}_{k}(t)=\text{MLP}_{\text{head}}\left([\mathbf{W}_{\text{self}}\mathbf{h}_{k}(t)\mathbin{\|}\mathbf{c}_{\text{neigh}}(t)]\right),$ (23)
@@
V-E Multi-Agent Actor-Critic Learning
@@
$L(\theta)=\mathbb{E}_{k,t}\left[\min(\rho_{k}(t)A_{k}(t),\text{clip}(\rho_{k}(t),1-\epsilon,1+\epsilon)A_{k}(t))\right],$ (24)
@@
$\mathcal{L}(\phi)=\begin{cases}\frac{1}{2}(y_{t}-V_{\phi}(s_{t}))^{2}&\text{for }|y_{t}-V_{\phi}(s_{t})|\leq\delta,\\ \delta(|y_{t}-V_{\phi}(s_{t})|-\frac{1}{2}\delta)&\text{otherwise},\end{cases}$ (25)
@@
V-F TAG-MAPPO Training Algorithm
The complete training procedure for the TAG-MAPPO framework is formalized in Algorithm 1. The optimization pipeline consists of decentralized data collection and centralized parameter updates.
@@ V-F TAG-MAPPO Training Algorithm
In the first phase, distributed UAV agents interact with the environment using lightweight actor networks. At each time step, agents rely on local observations to sample actions, ensuring inference latency remains constant regardless of network size. Upon node failures, the surviving swarm $\mathcal{V}(t)$ must immediately adapt based on remaining local observations. Transition data, including the instantaneous topology state $\mathcal{V}(t)$, is stored in the replay buffer to provide the critic with necessary context for post-failure value estimation.
@@
In the second phase, the system executes centralized updates. Before optimization, Generalized Advantage Estimation (GAE) is computed to balance the bias-variance trade-off. The TA-GAT critic plays a pivotal role by reconstructing dynamic ego-centric graphs for each minibatch sample. Specifically, the ROS mechanism is applied to the neighbor feature set before aggregation. By stochastically reordering neighbors, the critic learns relational features invariant to agent indexing. Through dual-path aggregation, the critic predicts state values to compute the Huber-based value loss. This mechanism ensures that lightweight actors become topology-aware by distilling complex relational knowledge from the critic into the simplistic policy networks.
V-G Computational Complexity Analysis
@@
During the online execution phase, computational complexity is dominated by matrix-vector multiplications within the MLP-based actors, scaling as $\mathcal{O}(L\cdot D^{2})$ where $L$ is the number of layers and $D$ is the hidden dimension. This complexity is independent of the number of dynamic neighbors, satisfying the real-time constraints required for 6G mission-critical tasks.
@@
During the offline training phase, complexity is primarily driven by the attention mechanism within the centralized critic. For each agent, the TA-GAT module involves linear projections and scaled dot-product computations. For an ego-centric graph, this scales as $\mathcal{O}((|\mathcal{V}(t)|+|\mathcal{N}_{k}(t)|)\cdot D)$, where $|\mathcal{N}_{k}(t)|$ is the number of active neighbors. While the attention operation involves more intensive processing than simple MLPs, its execution is restricted to the centralized training server with substantial resources. Consequently, TAG-MAPPO achieves an optimal trade-off by utilizing topology-aware reasoning offline to distill robust policies into efficient, lightweight actors for deployment.
VI Simulation Results and Analysis
@@
In this section, we conduct extensive simulations to evaluate the performance and resilience of the proposed TAG-MAPPO framework. We outline the simulation setup and analyze experimental results across various environmental scenarios, focusing on system recovery following node failures.
VI-A Simulation Setup and Parameters
@@
The simulation environment is developed using a customized Python-based multi-UAV simulator. The target area is a 1000×1000 $\times 1000\text{ m}^{2}$ square region served by a heterogeneous network consisting of $K_{U}=4$ mobile UAVs and one stationary ground base station (GBS), totaling $K=5$ serving nodes for $M=140$ users. The UAV flight altitude corridor is $H_{\min},H_{\max}]=[80,120]\text{ m}$. Maximum speeds for UAVs and independent ground users are capped at $V_{\max}^{\text{UAV}}=30\text{ m/s}$ and $V_{\max}^{\text{user}}=15\text{ m/s}$, respectively. In clustered scenarios, the group velocity for vehicular platoons is maintained at $10\text{ m/s}$.
@@
To ensure balanced learning across objectives, individual reward components in (14) are normalized to $[0,1]$. The weighting coefficients are $(\lambda_{\text{cov}},\lambda_{\text{ee}},\lambda_{\text{jr}},\lambda_{\text{min}})=(3.5,0.1,0.5,0.5)$, prioritizing global coverage while maintaining service quality for worst-case users. Additionally, the handoff penalty $\lambda_{\text{ho}}$ is tuned to $0.1$ to suppress redundant signaling and maintain stability during reconfiguration.
@@
To evaluate resilience, a catastrophic node failure is triggered at $t=100$ by setting $\xi_{k}(t)=0$ for a randomly selected UAV node (excluding the GBS), forcing the remaining swarm to reconfigure autonomously. Consolidated parameters are detailed in Table I.
VI-B Benchmark Scenarios and Channel Characterization
@@
To evaluate generalization and robustness, we define three simulation scenarios based on 3GPP propagation models. The specific LoS/NLoS characteristics for each environment are detailed as follows:
@@
  • Crowded Urban: This interference-limited regime features high user density with channel parameters $(a=12.08,b=0.11)$ and a significant NLoS path loss exponent ($\eta_{\text{nlos}}=23.0$). The primary challenge is managing complex co-channel interference within user clusters modeled via RPGM mobility.
@@
  • Suburban: This scenario involves moderate user dispersion and balanced LoS probability ($a=9.61,b=0.16$). It serves as the baseline for evaluating the optimal resilience of the UAV swarm, where the topology allows for effective spatial reuse.
@@
  • Rural Area: This noise-limited environment features sparse, Gauss-Markov independent mobility with high LoS probability ($a=4.88,b=0.43$). Lower blockage frequency simplifies link optimization but tests the over-smoothing limits of the graph coordination mechanism.
TABLE I: Consolidated Simulation Parameters and Hyperparameter Settings
Parameter (Physical) Symbol Value
Number of UAVs $K_{U}$ 4
Ground Base Station (GBS) - 1
UAV Max Speed $V_{\max}^{\text{UAV}}$ 30 m/s
User Max Speed $V_{\max}^{\text{user}}$ 15 m/s
Carrier Frequency $f_{c}$ 2.0 GHz
System Bandwidth $B$ 20 MHz
Transmission Power $P_{\text{tx}}$ 23 dBm
Noise PSD $N_{0}$ 174-174 dBm/Hz
Altitude Range $H_{\min},H_{\max}] [$80,120][80,120] m
GBS Fixed Power $P_{\text{GBS}}$ 20 W
UAV Comm. Power $P_{\text{comm}}$ 5 W
Target Area $\mathcal{D}$ 1×1 $\times 1\text{ km}^{2}$
Safety Distance $d_{\text{safe}}$ 10 m
Number of Users $M$ 140
Hyperparameter (DRL) Symbol Value
Actor Learning Rate $\alpha_{a}$ $1\times 10^{-4}$
Critic Learning Rate $\alpha_{c}$ $5\times 10^{-4}$
Discount Factor $\gamma$ 0.99
GAE Parameter $\tau$ 0.95
PPO Epochs $E_{\text{ppo}}$ 10
Batch Size $B_{\text{size}}$ 256
Huber Loss Delta $\delta$ 2.0
Entropy Coeff. $\beta_{\text{ent}}$ $0.10\to 0.01$
Total number of episodes $M$ 5000
Episode index epep 1, 2, …, 5000
Reward Weights Symbol Value
Coverage $\lambda_{\text{cov}}$ 3.5
Energy Efficiency $\lambda_{\text{ee}}$ 0.1
Rate Fairness $\lambda_{\text{jr}}$ 0.5
Minimum Rate $\lambda_{\text{min}}$ 0.5
Handoff Penalty $\lambda_{\text{ho}}$ 0.1
VI-C Performance Metrics